OpenStudy (anonymous):

given that f(x)=11/x-9 and g(x)=14/x+9 find f+g

OpenStudy (anonymous):

Find a common denominator so you can add them together.

OpenStudy (anonymous):

Is it \[\frac{ 11 }{ x-9 }\] or \[\frac{ 11 }{ x } -9\] Because that's a very important distinction

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

\[\frac{ 11(x+9) }{ (x-9) }+\frac{ 14(x-9) }{ (x+9) } = answer\]

OpenStudy (anonymous):

In that case, you need to find a common denominator, like hsrikako said. The easiest way to find a common denominator is to just multiply the two denominators together. BUT, we can't just do that with equations, so we have to multiply by 1. For example \[\frac{ 11 }{ x-9 }*\frac{ x+9 }{ x+9 }\] And then we multiply our second equation by \[\frac{ 14 }{ x+9 }*\frac{ x-9 }{ x-9 }\] This gives us a common denominator So then we can add them together to get: \[\frac{ 11*(x+9) + 14*(x-9) }{ (x-9)*(x+9) }\] if we multiply everything out, we get: \[\frac{ 11x+99 + 14x-126 }{ (x^2 -81) }\] which simplifies to \[\frac{ 25x - 27 }{ x^2 - 81 }\]

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

can you help me figure the f/g out